OpenStudy (anonymous):

Solve the following equations over [0°, 360°]. 1) 2 sin (x − 60)° = 1 2) √2 cos(x+90)°+1 = 0 Show working out thanks.

OpenStudy (callisto):

For question (1) 2 sin (x − 60)° = 1 Divide both sides by 2, what do you get?

OpenStudy (anonymous):

sin(x-60)°=1/2

OpenStudy (callisto):

Yes. Next, take arcsine for both sides, what do you get?

OpenStudy (anonymous):

sorry, what's arcsine?

OpenStudy (anonymous):

opposite of sine?

OpenStudy (callisto):

arcsine is the inverse of sine function.

OpenStudy (anonymous):

(x-60)°= sin-1 (1/2)

OpenStudy (anonymous):

sin^-1

OpenStudy (callisto):

Yes. What does \(sin^{-1}\frac{1}{2}\) give you?

OpenStudy (anonymous):

30

OpenStudy (callisto):

Yes. That's the angle in quadrant I, what about in quadrant IV?

OpenStudy (anonymous):

360-30=330 deg

OpenStudy (anonymous):

Or it's the good habit to memorize 1/2 = sin 30

OpenStudy (anonymous):

I do remember it actually.

OpenStudy (callisto):

Oh~ So.. you get x - 60 = 30 or x-60 = 330 Can you solve the two?

OpenStudy (anonymous):

x=90 or x=390

OpenStudy (callisto):

Ah!!! My mistake!! The other angle should be in quadrant II... Since sine is positive in quadrant I and II.. I'm sorry!!

OpenStudy (callisto):

I'm sorry... Just start it again! For question 1, 2 sin (x − 60)° = 1 Divide both sides by 2, which you've done just now. sin (x − 60)° = 1/2 Take arcsine for both sides, you get x-60 = 30 or x-60 = ______ (____is the angle in quadrant II, since sine of a angle in quadrant I and II gives you a positive value)

OpenStudy (anonymous):

kk

OpenStudy (callisto):

Can you work out the ___ ?

OpenStudy (anonymous):

Would it be 90?

OpenStudy (callisto):

Hmm.... We're still on x-60 = (angle in quadrant II), right?

OpenStudy (anonymous):

yeh but I'm a bit confused with your method.

OpenStudy (anonymous):

on that step.

OpenStudy (anonymous):

how I was taught to do it was to find the angle in the first quadrant first, in this case it is 90 I believe?

OpenStudy (anonymous):

and then using the angle that you found you use 180-that angle found= to get the answer

OpenStudy (callisto):

Can we do it in this way? Solve siny = 1/2, what is/ are y?

OpenStudy (anonymous):

y= sin^-1 (1/2)

OpenStudy (callisto):

which is equal to?

OpenStudy (anonymous):

30

OpenStudy (callisto):

and?

OpenStudy (anonymous):

150?

OpenStudy (callisto):

Right! So, you get y=30 or y=150, agree?

OpenStudy (anonymous):

yeh

OpenStudy (callisto):

Now, back you your question. 2 sin (x − 60)° = 1 sin (x-60) = 1/2 So, in this case, let x-60 = y, you'll get siny = 1/2. and just now, you solve y, which is 30 or 150. So can you get x?

OpenStudy (anonymous):

x-60=30 or x-60=150 x=90 or x= 210

OpenStudy (callisto):

Now, it's done~

OpenStudy (anonymous):

yeh but I'm still a bit confused with a part of it.

OpenStudy (callisto):

Which part?

OpenStudy (anonymous):

near the last part?

OpenStudy (callisto):

What is it exactly?

OpenStudy (anonymous):

"So, in this case, let x-60 = y, you'll get siny = 1/2. and just now, you solve y, which is 30 or 150. So can you get x?"

OpenStudy (anonymous):

wait, is it possible if u type out all the working out in logical order?

OpenStudy (callisto):

Oh... perhaps I should write it in this way. 2sin(x-60) = 1 Let y = x-60, then the equation becomes 2siny = 1 So, siny = 1/2 And as what you did, y= 30 or y=150 So, replace y by x-60 x-60 = 30 or x-60 = 150 Solve x. It should be clearer now. Please ask if you still have questions. And sorry for my poor presentation just now :(

OpenStudy (anonymous):

Or you can visualize the circle: 1/2 = sin30 = sin ( 180 - 30 )

OpenStudy (anonymous):

kk, I get the working out but yeh I may not remember how to do it your way as I was sort of taught a different way by my cousin and I absolutely hate these type of confusing type questions.

OpenStudy (callisto):

It's okay. Just choose the best way for yourself. :) Can you do question 2?

OpenStudy (anonymous):

yep, well I'll try do it your way and I think I can only do half of the steps

OpenStudy (callisto):

You can try to do it using your own way first :)

OpenStudy (callisto):

I'll check it for you

OpenStudy (anonymous):

wait, so I've done this so far using your method, the rest I don't get. √2 cos(x+90)°+1 = 0 cos(x+90)° = -1/√2 let y = x + 90 therefore cos y = -1/√2 y = ............ or y = ............

OpenStudy (callisto):

So far so good :) What are the values of y such that cosy = -1/√2

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

Hint: -1/√2 = -√2/ 2

OpenStudy (callisto):

Hmm.. Forget my method... Can you solve it using your OWN way??

OpenStudy (anonymous):

well it was my cousin's way and he only taught me just then. I still don't quite fully understand his method that's why I was planning to see if anyone could simplify it down for me on here.

OpenStudy (callisto):

Can you do me a favour by showing how your cousin has taught you? You can use another question or simply this question if you like.

OpenStudy (anonymous):

tan (x+45°) = 1 x∈[0°,360°] x+45 = 45° (x+45°) ∈ [45°,405°] 180° + 45° or 405° x+45° = 45°, 225° or 405° x = 0°, 180° or 360°

OpenStudy (anonymous):

Your cousin's way is exactly as same as our way!

OpenStudy (anonymous):

lol, really?

OpenStudy (callisto):

I hope you don't mind me asking the following question. 1. Have you learnt CAST ? 2. Do you know something like cos(180 - x) = -cosx sin(180-x) = sinx and so on? 3. Are you familiar with the trigo functions of special angles (that are 0 , 30 , 45, 60, 90, and so on) ? It's practically the same

OpenStudy (anonymous):

yes, I have and I do know the trig functions that u mentioned

OpenStudy (callisto):

Oh great. Sudden quiz. What is cos135?

OpenStudy (anonymous):

in exact value or degrees?

OpenStudy (anonymous):

Hint: 135 < 180

OpenStudy (callisto):

135 is in degrees. cos135 gives you an exact value.

OpenStudy (anonymous):

-cos45 = -Sqrt2/2

OpenStudy (callisto):

Yes.

OpenStudy (callisto):

Can we go back to your question 2 now?

OpenStudy (anonymous):

yes, please. It's late over here and I want to go to sleep lol.

OpenStudy (anonymous):

but I want to figure out how to do it properly first so I can do the other ones tomorrow.

OpenStudy (callisto):

Don't worry, I have to do my quiz in an hour too :| Here's your work! √2 cos(x+90)°+1 = 0 cos(x+90)° = -1/√2 let y = x + 90 therefore cos y = -1/√2 Can you solve y now?

OpenStudy (anonymous):

wait. let me see... and good look for your quiz.

OpenStudy (anonymous):

well since the value is negative it has to be in the second or third quadrant?

OpenStudy (callisto):

Yes. Another hint given by @Chlorophyll before : -1/√2 = - (1/√2) x (√2/√2) = - √2/2

OpenStudy (anonymous):

from there I'm a bit stuck as to how to solve for y.

OpenStudy (anonymous):

Consider y as X!

OpenStudy (callisto):

cosy = -√2/2 What is the value of angle such that cos(angle) = -√2/2? Another hint: refer to the sudden quiz

OpenStudy (anonymous):

yes, but how did you get the -√2/2?

OpenStudy (callisto):

Ratinoalization?! -1/√2 = - (1/√2) x (√2/√2) = - √2/2

OpenStudy (anonymous):

can't u cancel the top and bottom Sqrt 2 and so u get -1/Sqrt2

OpenStudy (callisto):

Oh yes, of course! But you need to know that cos45 = 1/√2

OpenStudy (anonymous):

-√2/2 looks more familiar to you or not?

OpenStudy (anonymous):

yes, it is cos 45 right?

OpenStudy (anonymous):

-cos 45 in this case*

OpenStudy (callisto):

-cos45 = cos(___ -45) = cos(____+45) ?

OpenStudy (anonymous):

-cos45 = cos(180-45) = cos (180+45)?

OpenStudy (callisto):

Yes. So... cosy = -1/√2 y = ... or y= ...?

OpenStudy (anonymous):

wait... let me write something else down first.

OpenStudy (callisto):

Take your time~

OpenStudy (anonymous):

@JayDS @Callisto giant leap improvement here =)

OpenStudy (anonymous):

lol, no way, I haven't improved at all but Callisto has.

OpenStudy (callisto):

What??!!! You haven't improved at all :'( ???

OpenStudy (anonymous):

√2 cos(x+90)°+1 = 0 cos(x+90)° = -1/√2 let y = x + 90 therefore cos y = -1/√2 cos y = - √2/2 hmm where to from here?

OpenStudy (anonymous):

I guess I have slightly lol, I'm just really stupid.

OpenStudy (anonymous):

@JayDS Our goal here is to help YOU improved :)

OpenStudy (callisto):

No one is stupid except me, you can't be stupid than me:| √2 cos(x+90)°+1 = 0 cos(x+90)° = -1/√2 let y = x + 90 So, the equation becomes cos y = -1/√2 Your duty: solve y. How...?!